CASE STUDY
Baseball Card Emporium
Baseball Card Emporium (BBE) of Lewistown, Pennsylvania, is a distributor of baseball
cards to sports card retailers. Its market area encompasses most of Pennsylvania, eastern
Ohio, and New Jersey. The cards are printed in Neenah, Wisconsin, and currently shipped
to Lewistown via motor carrier transportation. Kenny Craig, vice president of logistics, has
asked his staff to evaluate using air carrier service to ship the cards.
Nick Gingher, director of distribution, has collected the following information:
• Annual demand: 6,000 cases of cards
• Case value (price): $96 each
• Inventory carrying cost (annual): 30%
• Cost per order to replenish inventory: $75
• In-transit inventory carrying cost: 18%
• Transit time using motor carrier (parcel ground: 4 days
• Cycle time using motor carrier (parcel ground): 7 days
• Transit time using air carrier: 1 day
• Cycle time using air carrier:2 days
• Motor carrier rate: $1.20 per cwt. (100 lbs.)
• Air carrier rate: $2.50 per cwt.
• Unit weight: 50 lbs. per case
CASE QUESTIONS
1. What is the economic order quantity for BBE in units? In pounds?
2. What is the total cost (not considering transportation-related costs) of the EOQ?
3. What is the total cost for using motor carrier transportation?
4. What is the total cost for using air carrier transportation?
5. Which alternative should BBE use?
Answer #1

1. EOQ = [ 2 xdemand per annum x ordering cost / holding cost per unit per year]1/2

= [ 2x 6000x75/96x0.3]1/2 = 176.77 =177

2. TC of EOQ

= Purchase cost + holding cost + ordering cost

= 96x6000+ 177/2 x 96x0.3 + 6000/177 x 75 = 581091

3. TC using motor carrier transportation

= TC of EOQ + carrier cost + holding cost

= 581091 + 1.2 x6000/20+ 4/365 x0.18x96 x6000 = 582587

4.TC using air carrier transportation

= 581091 + 2.5 x6000/20 + 1/365x0.18x96x6000 = 582185

The air carrier option is more cost effective.